# Compress then Serve: Serving Thousands of LoRA Adapters with Little Overhead

### v2

@@ Appendix D Avoiding Batched Matrix Multiplication (BMM)
Appendix E GPU Memory Usage Computation for JD Compression
@@
  • $N$: Maximum number of LoRA modules being served simultaneously (max_lora_num).
@@
In Figure 1, we use different JD-compression settings for serving different number of unique LoRAs. SpecificallyL
  • Serving 4 unique LoRAs. Ours: rank 16 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 2
  • Serving 8 unique LoRAs. Ours: rank 16 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 2
  • Serving 16 unique LoRAs. Ours: rank 32 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 3
  • Serving 32 unique LoRAs. Ours: rank 64 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 5
  • Serving 64 unique LoRAs. Ours: rank 64 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 6
  • Serving 128 unique LoRAs. Ours: 7 clusters rank 16 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 8
  • Serving 256 unique LoRAs. Ours: 10 clusters rank 16 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 10
  • Serving 512 unique LoRAs. Ours: 25 clusters rank 16 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 26
  • Serving 1024 unique LoRAs. Ours: 7 clusters rank 64 JD-Full. vLLM multiLoRA baseline: max-gpu-lora is 60
E.1 Baseline GPU Memory Usage
@@
E.2 GPU Memory Usage for JD Full Method
@@
E.3 GPU Memory Usage for Clustering Method
@@
Appendix F Selecting Number of Clusters
To identify optimal hyperparameters for the clusters compression method, we analyzed the relationship between reconstruction error and the parameter saved ratio for a single LoRA module, as shown in Figure 5. By comparing the results across different numbers of Low-Rank Adaptation (LoRA) configurations (100 and 500, depicted in subfigures 5(a) and 5(b)), we were able to observe the trade-off between model size reduction and reconstruction accuracy. Based on these findings, we selected the rank and number-of-clusters hyperparameters that effectively balance these two objectives. The chosen settings were then used to conduct full-scale experiments.
@@ Appendix F Selecting Number of Clusters
Appendix G Additional Results
@@
G.1 Relative Rouge-L Performance and Compression Rate

### v3

@@ Appendix D Avoiding Batched Matrix Multiplication (BMM)
Appendix E Simple Timing Experiments
In Figure 5, we present a set of simple experiments comparing the memory load, transfer time, and forward-pass performance of LoRA and JD-LoRA. These experiments were conducted across multiple clusters and various rank configurations. For memory usage, we measured all 96 LoRA modules in the Mistral model; however, for transfer time and forward-pass performance, we tested only a single LoRA module. We also implemented F-LoRA (Wen & Chaudhuri, 2024), but contrary to their reported results, we were unable to achieve faster forward-pass performance than standard LoRA.
Figure 5: Memory load, transfer time, and forward-pass performance of LoRA and JD-LoRA.
Appendix F GPU Memory Usage Computation for JD Compression.
@@ Appendix E Simple Timing Experiments
  • NN: Maximum number of LoRA modules being served simultaneously (max_lora_num).
@@
In Figure 1, we use different JD-compression settings for serving different number of unique LoRAs. Specifically:
  • Serving 4 unique LoRAs:
     Ours: rank 16 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 2.
  • Serving 8 unique LoRAs:
     Ours: rank 16 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 2.
  • Serving 16 unique LoRAs:
     Ours: rank 32 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 3.
  • Serving 32 unique LoRAs:
     Ours: rank 64 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 5.
  • Serving 64 unique LoRAs:
     Ours: rank 64 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 6.
  • Serving 128 unique LoRAs:
     Ours: 7 clusters, rank 16 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 8.
  • Serving 256 unique LoRAs:
     Ours: 10 clusters, rank 16 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 10.
  • Serving 512 unique LoRAs:
     Ours: 25 clusters, rank 16 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 26.
  • Serving 1024 unique LoRAs:
     Ours: 25 clusters, rank 16 JD-Full.
     vLLM multiLoRA baseline: max-gpu-lora = 28
F.1 Baseline GPU Memory Usage
@@
F.2 GPU Memory Usage for JD Full Method
@@
F.3 GPU Memory Usage for Clustering Method
@@
F.4 Punica
In our vLLM experiments, we specifically used the Punica kernel for implementing multi-LoRA, applying our approach in conjunction with Punica’s capabilities. Our custom function, add_lora_slice_with_sigma, implements the following key steps:
  1. 1.
    Initialize Buffers: Creates temporary storage for intermediate calculations if not already provided.
  2. 2.
    Apply Matrix A: Transforms x using matrix A, storing the result in buffer.
  3. 3.
    Apply Matrix Sigma: Further transforms buffer using Sigma, storing the result in buffer_sigma.
  4. 4.
    Apply Matrix B and Update y: Finally, transforms buffer_sigma using B, applies scaling, and updates a slice of y in place.
Below is the pseudocode for add_lora_slice_with_sigma, illustrating the integration:
Listing 1: Pseudocode for ‘add_lora_slice_with_sigma‘
1Function add_lora_slice_with_sigma(y, x, wa_t_all, wb_t_all, wsigma_t_all, indices, layer_idx, scale, y_offset, y_slice_size, buffer=None):
2 # Initialize buffers if not provided
3 if buffer is None:
4 buffer = create_tensor(shape=(x.size(0), R), dtype=float32)
5 buffer_sigma = create_tensor(shape=(buffer.size(0), R), dtype=float32)
6 # Step 1: Apply matrix A
7 dispatch_bgmv_low_level(buffer, x, wa_t_all, indices, layer_idx, scale=1.0)
8 # Step 2: Apply matrix Sigma
9 dispatch_bgmv_low_level(buffer_sigma, buffer, wsigma_t_all, indices, layer_idx, scale=1.0)
10 # Step 3: Apply matrix B and update y slice
11 dispatch_bgmv_low_level(y, buffer_sigma, wb_t_all, indices, layer_idx, scale, y_offset, y_slice_size)
12End Function
Appendix G Selecting Number of Clusters
To identify optimal hyperparameters for the clusters compression method, we analyzed the relationship between reconstruction error and the parameter saved ratio for a single LoRA module, as shown in Figure 6. By comparing the results across different numbers of Low-Rank Adaptation (LoRA) configurations (100 and 500, depicted in subfigures 6(a) and 6(b)), we were able to observe the trade-off between model size reduction and reconstruction accuracy. Based on these findings, we selected the rank and number-of-clusters hyperparameters that effectively balance these two objectives. The chosen settings were then used to conduct full-scale experiments.
@@ Appendix G Selecting Number of Clusters
Appendix H Additional Results
@@
H.1 LoRAs of different ranks
In Table 4, we report the performance of our compression method on LoRAs with ranks uniformly sampled between 16 and 64 (mean rank of 43). In Table 5, we present compression results for LoRAs of rank 43, matching the average rank from Table 4. Because these same-rank LoRAs have an identical parameter count, they also exhibit identical parameter-saving ratios. While performance declines slightly for LoRAs with varying ranks, our compression method still preserves over 99% of the original performance.
Table 4: Results for Different Numbers of Clusters (different LoRA-ranks)
Metric Uncompressed 1 Cluster 2 Clusters 4 Clusters 8 Clusters
Loss 0.5009 1.6023 0.9986 0.5603 0.5000
Exact match 69.6103 37.3392 60.1071 66.6474 69.6289
ROUGE-1 79.5755 50.3598 71.5644 76.2930 79.0576
ROUGE-L 78.9355 49.7491 71.0193 75.6152 78.3958
Recon. error 0 0.8311 0.6990 0.4846 0.3246
Agreement 1.0 36.6706 62.6699 72.2231 80.5545
Exact match ratio 1.0 0.5724 0.8079 0.9219 0.9541
Relative ROUGE-1 1.0 0.6220 0.8533 0.9406 0.9917
Relative ROUGE-L 1.0 0.6221 0.8550 0.9404 0.9915
Param. saved ratio 1.0 0.99 0.98 0.97 0.94
Table 5: Results for Different Numbers of Clusters (same LoRA-ranks of 43)
Metric Uncompressed 1 Cluster 2 Clusters 4 Clusters 8 Clusters
Loss 0.5764 0.7767 0.6277 0.5841 0.5659
Exact match 61.7746 53.2000 60.2000 61.5000 61.3000
ROUGE-1 79.9322 74.7695 79.4621 80.6950 80.4298
ROUGE-L 77.7961 72.4257 77.2141 78.3369 78.1883
Recon. error 0 0.7333 0.5594 0.3999 0.2502
Agreement 1.0 0.0000 6.6667 6.6667 6.6667
Exact match ratio 1.0 0.8175 0.9261 0.9618 1.0206
Relative ROUGE-1 1.0 0.9560 1.0153 1.0310 1.0241
Relative ROUGE-L 1.0 0.9501 1.0118 1.0268 1.0231
Param. saved ratio 1.0 0.99 0.98 0.97 0.94
H.2 Privacy Ablation
We investigate whether jointly compressing certain tasks results in improved performance on tasks within that same compressed group, compared to tasks outside of the group. In other words, we examine whether information about which tasks were compressed together could be inferred from subsequent performance differences. Table LABEL:table:privacy presents the results.
In the Compressed Together setting, tasks (task1391, task190, task280, task290, and task391) are compressed jointly with five other tasks. We then evaluate their cross-task performance within this group. In the Compressed Separately setting, the same set of tasks is each compressed alongside nine other tasks that are not part of the original group. We again evaluate the cross-task performance using the same sets of tasks, allowing us to compare and assess any differences in performance attributable to joint versus separate compression.
Table 6: Privacy Ablation: Performance on tasks compressed together vs. performance on tasks compressed separately
Source Task Task Combination Test Loss Exact Match Rouge1 RougeL
Compressed Together
task1391 task1391 on task190 2.320 29 29.00 29.00
task1391 on task280 1.445 10 12.00 12.00
task1391 on task290 2.662 0 0.00 0.00
task1391 on task391 1.560 4 7.00 7.00
Average 1.997 10.8 12.00 12.00
task190 task190 on task1391 1.392 0 0.00 0.00
task190 on task280 1.647 2 2.00 2.00
task190 on task290 1.838 0 0.00 0.00
task190 on task391 2.622 2 2.00 2.00
Average 1.875 1.0 1.00 1.00
task280 task280 on task1391 2.259 14 16.47 16.47
task280 on task190 2.922 20 20.50 20.50
task280 on task290 0.729 33 43.07 43.07
task280 on task391 2.318 44 62.52 62.52
Average 2.057 27.8 35.64 35.64
task290 task290 on task1391 0.867 36 46.58 46.58
task290 on task190 1.553 36 36.00 36.00
task290 on task280 1.036 43 43.40 43.40
task290 on task391 0.461 59 86.33 86.33
Average 0.979 43.5 53.08 53.08
task391 task391 on task1391 0.502 65 65.75 65.75
task391 on task190 1.421 31 31.00 31.00
task391 on task280 0.417 69 69.00 69.00
task391 on task290 0.265 71 90.33 90.33
Average 0.651 59.0 64.02 64.02
Compressed Separately
task1391 task1391 on task190 2.347 30 30.00 30.00
task1391 on task280 1.543 11 13.58 13.58
task1391 on task290 2.477 0 0.00 0.00
task1391 on task391 1.253 10 18.00 18.00
Average 1.905 12.8 15.40 15.40
task190 task190 on task1391 1.388 0 0.00 0.00
task190 on task280 1.603 2 2.00 2.00
task190 on task290 1.771 0 0.00 0.00
task190 on task391 2.326 4 4.00 4.00
Average 1.772 1.5 1.50 1.50
task280 task280 on task1391 3.111 2 5.13 5.13
task280 on task190 2.711 19 19.00 19.00
task280 on task290 0.948 16 22.09 22.09
task280 on task391 2.449 39 53.11 53.11
Average 2.305 19.0 24.83 24.83
task290 task290 on task1391 0.848 41 47.58 47.58
task290 on task190 1.355 38 38.00 38.00
task290 on task280 1.050 41 41.00 41.00
task290 on task391 0.463 59 86.33 86.33
Average 0.929 44.8 53.23 53.23
task391 task391 on task1391 0.428 68 68.74 68.74
task391 on task190 1.507 31 31.00 31.00
task391 on task280 0.368 70 70.00 70.00
task391 on task290 0.269 73 91.00 91.00
Average 0.643 60.5 65.18 65.18
Table 6: Privacy Ablation: Performance on tasks compressed together vs. performance on tasks compressed separately (continued)
H.3 Relative Rouge-L Performance and Compression Rate